OpenStudy (anonymous):

Help meeee. What is the first step in solving this system of equations using elimination? y=3/4x+3 and 2x-4y=6

OpenStudy (anonymous):

why not use substitution, since you already know \(y=\frac{3}{4}x+3\)

OpenStudy (anonymous):

But thats not what the question is asking me );

OpenStudy (anonymous):

then i guess the first step would be to write it as \[\frac{3}{4}x-y=-3\]

OpenStudy (anonymous):

so you would have the two equations \[2x-4y=6\\ \frac{3}{4}x-y=-3\]

OpenStudy (anonymous):

Thank you.